A roller coaster cart with a mass of 1283 kg starts on a hill that is 189 m off the ground. The hill descends all the way to the ground, before going back up hill. How fast is the roller coaster going when it reaches the ground?
1
Expert's answer
2020-03-30T08:06:02-0400

From the conservation of energy:


"0.5mv^2=mgh"

"v=\sqrt{2gh}=\sqrt{2(9.81)(189)}=60.9\frac{m}{s}"